OpenStudy (keana):

Find the height of a square pyramid that has a volume of 12 cubic feet and a base length of 3 feet. 1.5 feet 2 feet 3 feet 4 feet

OpenStudy (hoblos):

the volume of a square pyramid is \[V = \frac{ s^2 h }{ 3 } \] here v = 12 and s=3 solve for h

OpenStudy (keana):

what is s?

OpenStudy (keana):

is it the length of the side?

OpenStudy (hoblos):

the side of the base

OpenStudy (hoblos):

yes it is

OpenStudy (keana):

well how do I know what it is??

OpenStudy (hoblos):

replace s by 3 and v by 12 what do you get ?

OpenStudy (keana):

12=3^2h/3

OpenStudy (keana):

12=9h/3

OpenStudy (keana):

so is it 3?

OpenStudy (hoblos):

correct so now we have to solve for h first multiply both sides by 3 what do you get ?

OpenStudy (keana):

which sides??

OpenStudy (hoblos):

sides of the equation it means what before the = and what after it